Problem 10:

Browny Brad’s Beach Motel (BBBM), a forty-room lodging facility, has operated for the past three years. BBBM’s night auditor has kept accurate records over the past year but has not analyzed any of these data. The following is a summary of the rooms sold by month:

Rooms sold:

            January      400                        March         700                          May           960

            February   600                         April            840                         June            980

            July.           992                        September   800                       November   650

            August.     973                         October     705                        December    500

Assume that the motel is open 365 days per year for business purposes and that the motel considers its “summer months” to be May, June, July, and August. The off-season comprises of all other months.

Required:

Determine the Paid Occupancy percentage for the summer months, the off-season, and the entire year.
Answer #1

Total Room Nights available:

May-40*31=1240

June-40*30=1200

July=40*31=1240

August=40*31=1240

Total Room nights available in Summer months =1240+1200+1240+1240=4920

Total Occupancy in summer months=960+980+992+973=3905

Paid Occupancy percentage for the summer months=(3905/4920)*100=79.37%

Room Nights available during entire year =365*40=14600

Room Nights available during off season=14600-3905=10695

Total Occupancy in Off Season:

January;400

February:600

March:700

April:840

September:800

October: 705

November:650

December:500

Total Occupancy in off season=400+600+700+840+800+705+650+500=5195

Paid Occupancy percentage for the off season=(5195/10695)*100=48.57%

Total Occupancy in the whole year=3905+5195=9100

Room Nights available during entire year =14600

Paid Occupancy percentage for the entire year=(9100/14600)*100=62.33%